### v1

@@ Abstract
We develop a framework for the dynamical closure of spatiotemporal dynamics governed by partial differential equations. We employ the mathematical framework of quantum mechanics to embed the original classical dynamics into an infinite dimensional quantum mechanical system, using the space of quantum states to model the unresolved degrees of freedom of the original dynamics and the technology of quantum measurement to predict their contributions to the resolved dynamics. We use a positivity preserving discretization to project the embedded dynamics to finite dimension. Combining methods from operator valued kernels and delay embedding, we derive a compressed representation of the dynamics that is invariant under the spatial symmetries of the original dynamics. We develop a data driven formulation of the scheme that can be realized numerically and apply it to a dynamical closure problem for the shallow water equations, demonstrating that our closure model can accurately predict the main features of the true dynamics, including for out of sample initial conditions.
@@
dynamical closure, parametrization, quantum mechanics, kernel methods, delay embedding, transfer operators, shallow water equations.
@@ 1 Introduction
Complex dynamical systems often encompass degrees of freedom that evolve over a wide range of temporal and spatial scales, with their evolution depending on different physical regimes. A challenging problem in such multiscale, multiphysics systems is trying to include dynamical information for degrees of freedom that either are too computationally expensive to simulate directly or depend on fully or partially unknown physical processes [14].
The approach employed by closure or parametrization methods is to derive surrogate dynamical models that can be used to approximate the contributions of fine grain degrees of freedom to the dynamics of coarse grain processes we can simulate consistently. Such approaches have their origin in climate dynamics and turbulence research [38, 33, 1, 18], but have since found consequential applications in a wide range of fields [15].
Mathematically, closure methods depend on a separation of the state space of the original dynamical system into two components: (1) the resolved degrees of freedom, which represent the coarse grain processes we can simulate or measure consistently and with tractable cost; (2) the unresolved degrees of freedom, which are considered too complex or expensive to simulate or measure directly. Once the unresolved degrees of freedom have been identified, one has to define a surrogate model for their fluxes—namely, their contributions—to the resolved dynamics. How to define such a model, and whether such a model should be deterministic or stochastic, are active areas of research that use various mathematical tools to offer alternative approaches to this problem [14, 29, 2].
One approach toward this problem is defining a surrogate model as a functional relationship between the resolved and unresolved degrees of freedom. In this way, the unresolved variables are determined by a deterministic function of the resolved variables. Early attempts to define such functions depended on expert knowledge of the underlying physical process being modeled, featuring formulas that try to predict the averaged fluxes to the resolved variables. More recently, an alternative data driven approach has emerged, where functional relationships are not derived based on expert physical reasoning but are discovered by analyzing large amounts of numerical or experimental data [5, 7, 8, 24, 28, 30, 32, 44].
@@
To address the issue of dimension reduction and the potential loss of complex dynamics, one approach has been to use stochastic methods to model the unresolved degrees of freedom, instead of deterministic ones [10, 2, 29]. In this way, the surrogate unresolved variables retain some of their dynamical independence from the resolved variables, and the state space of the parametrized system is not reduced to the resolved state space.
@@
Independently of whether the unresolved variables are represented using a deterministic or stochastic framework, the main goal of closure methods is to construct a surrogate model that can accurately and consistently approximate the original dynamics. Additionally, it should do so with a tractable computational cost and without imposing a strong reduction in the complexity of the predicted dynamical behavior.
In the present work we extend the framework developed in [20, 16, 17] to the closure of spatiotemporal dynamics, modeling the unresolved dynamics using a quantum mechanical system. More specifically, the unresolved degrees of freedom are represented by quantum density operators and the fluxes by quantum observables, while the generation of the required flux terms is carried out using tools from the mathematical theory of quantum measurement [23, 34, 36]. The closure framework can be used to derive both deterministic and stochastic surrogate models; in this work we focus on the deterministic setting. In addition to structure preservation properties [16, 17], our proposed scheme can be realized numerically in a data driven fashion using kernel integral operator methods [3, 19] and spectral analysis of vector valued observables [22]. Relying on operator theoretic methods in Hilbert spaces, the data driven framework has the additional benefit of a rigorous consistency proof in the limit of infinite training data under reasonable assumptions [16, 17].
1.1 Problem statement
@@
We consider a dynamical system $\Phi^{t}:X\rightarrow X$, $t\geq 0$, on an infinite dimensional state space $X$ possessing an ergodic invariant probability measure $\mu$ with compact support. The flow map $\Phi^{t}$ represents the spatiotemporal dynamics governed by a partial differential equation on a spatial domain $S$. As such, elements of state space $X$ are functions on the spatial domain $S$; more specifically, we consider $X\subset L^{2}(S,\nu;\mathbb{R})$ with $\nu$ denoting the Lebesgue measure on $S$.
@@
We assume that there exists a symmetry group $G^{\prime}$ acting on the spatial domain $S$ with a continuous left action $\Gamma_{S,g}:S\rightarrow S$, $g\in G^{\prime}$ which preserves the null sets of measure $\nu$. The action $\Gamma_{S,g}$ induces the action on state space $\Gamma_{X,g}:X\rightarrow X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$, $g\in G^{\prime}$. The actions of group $G^{\prime}$ on $X$ represent dynamical symmetries of the dynamics $\Phi^{t}$; namely, we assume the following equivariance property
@@
for all $t\geq 0$, $g\in G^{\prime}$.
We decompose the state space $X=\mathcal{X}\times\mathcal{Y}$ with $\mathcal{X}$ denoting the state space of the resolved degrees of freedom of the dynamics and $\mathcal{Y}$ the state space of the unresolved degrees. In the present work we operate under the assumption that the resolved state space $\mathcal{X}$ is a finite dimensional differentiable manifold that can be embedded into $\mathbb{R}^{d^{\prime}}$ for some $d^{\prime}\in\mathbb{N}$. In many applications it is natural to assume that $\mathcal{X}$ is finite dimensional. For instance many dissipative partial differential equations give rise to long time dynamics that is confined to a finite dimensional subset of their state space [37]. In other cases discretization of the differential operator modeling the considered dynamics also yields a finite dimensional state space.
We introduce a flux space $\mathcal{Z}=\mathbb{R}^{d}$, $d\in\mathbb{N}$, and flux map $Z:\mathcal{Y}\rightarrow\mathcal{Z}$ which will allow us to include in the resolved dynamics terms that depend on the unresolved variables. Based on our decomposition of state space $X$ we define the flow maps for the resolved dynamics $\Phi_{\mathcal{X}}^{t}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{X}$, $\Phi_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\Phi^{t}$ and unresolved dynamics $\Phi_{\mathcal{Y}}^{t}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{Y}$, $\Phi_{\mathcal{Y}}^{t}=\operatorname{proj}_{\mathcal{Y}}\circ\Phi^{t}$, where $\operatorname{proj}_{A}$ denotes the canonical projection onto a subspace $A$. With these definitions in place, the dynamics $\Phi^{t}$ can be expressed as
$\Phi^{t}(\hat{x},y)=\bigl{(}\Phi^{t}_{\mathcal{X}}(\hat{x},Z(y)),\Phi_{\mathcal{Y}}^{t}(\hat{x},y)\bigr{)}$
with state vector $(\hat{x},y)\in\mathcal{X}\times\mathcal{Y}$.
Our goal in this work is to construct surrogate dynamics $\tilde{\Phi}^{t}:\tilde{X}\rightarrow\tilde{X}$ in a surrogate state space $\tilde{X}=\mathcal{X}\times\tilde{\mathcal{Y}}$ where $\tilde{\mathcal{Y}}$ denotes a surrogate unresolved state space. Based on our distinction between $\mathcal{Y}$ and $\tilde{\mathcal{Y}}$ we also introduce a surrogate version of the flux map $\tilde{Z}:\tilde{\mathcal{Y}}\rightarrow\mathcal{Z}$. The surrogate dynamics $\tilde{\Phi}^{t}$ can now be expressed as
(1) $\tilde{\Phi}^{t}(\hat{x},\tilde{y})=\bigl{(}\tilde{\Phi}_{\mathcal{X}}^{t}(\hat{x},\tilde{Z}(\tilde{y})),\tilde{\Phi}_{\mathcal{Y}}^{t}(\hat{x},\tilde{y})\bigr{)}$
with state vector $(\hat{x},\tilde{y})\in\mathcal{X}\times\tilde{\mathcal{Y}}$, surrogate resolved dynamics $\tilde{\Phi}_{\mathcal{X}}^{t}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{X}$, $\tilde{\Phi}_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\tilde{\Phi}$ and surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{t}:\mathcal{X}\times\tilde{\mathcal{Y}}\rightarrow\tilde{\mathcal{Y}}$, $\tilde{\Phi}_{\mathcal{Y}}^{t}=\operatorname{proj}_{\tilde{\mathcal{Y}}}\circ\tilde{\Phi}$. To build the surrogate dynamics map $\tilde{\Phi}^{t}$ we need to define the surrogate space $\tilde{\mathcal{Y}}$, the surrogate flux map $\tilde{Z}$, and the surrogate dynamics maps $\tilde{\Phi}_{\mathcal{X}}^{t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{t}$.
1.2 Main results
In this work we extend the original quantum mechanical closure framework (QMCL) to the case of spatiotemporal dynamics governed by partial differential equations. The QMCL framework models the unresolved variables of a dynamical system as a quantum mechanical system, using notions from quantum mechanics to define the surrogate flux map and surrogate dynamics [20, 16, 17]. We combine the original QMCL scheme with elements of the vector valued spectral analysis (VSA) framework introduced in [22] to develop a scheme that can effectively and efficiently model the closure of spatiotemporal dynamics. The main features of the present work can be summarized as follows.
  1. 1.
    Positivity preservation. As with the original QMCL work, our framework uses the noncommutative setting of operator algebras to encode the original classical dynamics, leading to a closure scheme that is positivity preserving. This means that positive flux terms in the original dynamics are represented by positive operators in our closure model, producing positive surrogate flux terms. Generating surrogate fluxes that preserve the sign of the original fluxes is a nontrivial task and has important consequences for the closure of dynamics with sign definite quantities [17].
  2. 2.
    Symmetry factorization. Taking advantage of the symmetry factorization properties of the VSA framework, we derive a dynamical closure scheme that has important invariance properties under the actions of the spatial symmetry group of the original resolved dynamics. This enables us to form an especially compressed representation of the original spatiotemporal dynamics, leading to a data driven realization of our closure scheme that can be applied to complex dynamics with tractable computational cost [22].
  3. 3.
    Computational efficiency. In addition to the factorization of spatial symmetries, we implement the data driven formulation of our scheme in a way that lowers the cost of its most computationally expensive operations. As with the original QMCL and VSA frameworks, our data driven scheme requires solving an eigenvalue problem for a kernel matrix. In the setting of spatiotemporal dynamics, the dimensions of this matrix can be very large, making the eigenvalue problem one of the most expensive tasks in our scheme. To address that we perform a positivity preserving approximation of the kernel matrix by a lower rank matrix, and use the lower rank approximation to solve the eigenvalue problem with reduced cost [9, 40].
  4. 4.
    Application. We apply our scheme to a closure problem for the shallow water equations on a one dimensional spatial domain. The shallow water equations are a hyperbolic system of partial differential equations with important applications in fields such as water waves modeling and ocean dynamics [41, 42]. We demonstrate that our scheme can effectively capture and reproduce the main features of the dynamics, including for out of sample initial conditions.
1.3 Paper layout
In Section 2 we provide a brief overview of the original QMCL framework. In Section 3 we extend this framework to the closure of spatiotemporal dynamics. We present the main features of the extended scheme, the elements it borrows from the VSA framework, as well as its data driven realization. Next, in Section 4 we apply the extended scheme to a closure problem for the shallow water equations and present our numerical results. We conclude with a discussion of some aspects of the presented scheme and future work directions in Section 5. Appendix A provides additional details on the computational implementation of the data driven formulation of our closure scheme.

### v2

@@ Abstract
We develop a framework for the dynamical closure of spatiotemporal dynamics governed by partial differential equations. Employing the mathematical framework of quantum mechanics to embed the original classical dynamics into an infinite dimensional quantum mechanical system, we use the space of quantum states to model the unresolved degrees of freedom of the original dynamics and the framework of quantum measurement to predict their contributions to the resolved dynamics. The embedded dynamics is projected to finite dimension by a positivity preserving discretization process. Based on methods from operator valued kernels and delay embedding, the compressed finite dimensional representation of the dynamics is invariant under the spatial symmetries of the original dynamics. We develop a data driven formulation of the scheme that can be realized numerically and apply it to a dynamical closure problem for the shallow water equations. The results demonstrate that our closure model can accurately predict the main features of the true dynamics, including for out of sample initial conditions.
@@
dynamical closure, parametrization, quantum mechanics, kernel methods, delay embedding, transfer operator, shallow water equations
[…]
[1]organization=Department of Mathematics, Dartmouth College, city=Hanover, state=NH 03755, country=USA
@@ 1 Introduction
Complex dynamical systems often encompass degrees of freedom that evolve over a wide range of temporal and spatial scales with their evolution depending on different physical regimes. A challenging problem in such multiscale, multiphysics systems is trying to include dynamical information for degrees of freedom that either are too computationally expensive to simulate directly or depend on fully or partially unknown physical processes [1].
The approach employed by closure or parametrization methods is to derive surrogate dynamical models that can be used to approximate the contributions of fine grain degrees of freedom to the dynamics of coarse grain processes we can simulate consistently. Early applications include turbulent flows and climate dynamics [2, 3, 4, 5], but closure methods have since found consequential applications in a wide range of fields [6]. For example, in atmospheric dynamics a technique known as super-parametrization employs deterministic or stochastic column models with explicitly resolved cloud microphysics as surrogate models of small scale moist convection, embedded within the discretization cells of a coarse atmospheric model [7, 8]. In plasma dynamics there have been efforts to derive models for subgrid scale processes in fluid models as well as to derive fluid closures of kinetic models, incorporating both physical modeling and data driven approaches [9, 10, 11, 12, 13, 14, 15, 16].
Related to closure is model reduction, which aims to derive models of reduced complexity that agree in a prescribed sense with the original dynamics. For instance, reduced models can be derived using operator theoretic methods that attempt to justify the model’s physical consistency in a statistical sense [17, 18, 19, 20, 21] or directly through computational methods [22, 23, 24, 25, 26, 27, 28, 29].
Mathematically, closure methods depend on a separation of the state space of the original dynamical system into two components: (1) the resolved degrees of freedom, which represent the coarse grain processes we can simulate or measure consistently and with tractable cost; (2) the unresolved degrees of freedom, which are considered too complex or expensive to simulate or measure directly. Once the unresolved degrees of freedom have been identified, one has to define a surrogate model for their fluxes—namely, their contributions—to the resolved dynamics. How to define such a model and whether such a model should be deterministic or stochastic are active areas of research that use various mathematical tools to offer alternative approaches [30, 31, 1, 32, 33].
One approach toward this problem is defining a surrogate model as a functional relationship between the resolved and unresolved degrees of freedom. In this way, the unresolved variables are determined by a deterministic function of the resolved variables. A widespread approach of defining such functions that was especially popular early on relies on expert knowledge of the modeled physical process to derive formulas that seek to predict the averaged fluxes to the resolved variables [34]. These formulas typically feature tunable parameters that are calibrated by analyzing numerical, observational or experimental data through regression or inference techniques [35, 36, 37, 38]. More recently, an alternative data driven approach has emerged where functional relationships are not derived based on expert physical reasoning but are discovered from data using machine learning methods [39, 40, 41, 42, 43, 27, 44, 45, 28, 46, 47, 48, 49].
@@
To prevent the reduction of the full state space to the resolved state space one can introduce a surrogate state space for the unresolved degrees of freedom and use that space to model their contributions to the resolved dynamics. One such approach has been to use stochastic methods to model the unresolved degrees of freedom, thereby complementing the resolved state space with the event space of the introduced stochastic variables [51, 52, 31, 53, 32, 54].
@@
Independently of whether the unresolved variables are represented using a deterministic or stochastic framework, the main goal of closure methods is to construct a surrogate model that can accurately and consistently approximate the original dynamics. Additionally, it should do so with a tractable computational cost and without imposing a strong reduction in the complexity of the predicted dynamical behavior. Despite the wide variety of mathematical tools employed, reduction and closure methods can be studied within a unified operator theoretic framework. Often referred to as the Mori-Zwanzig formalism, the framework uses projection operators to describe the reduction of fine grain degrees of freedom to coarse grain ones, resulting in exact governing equations for the coarse grain variables [17, 18, 19, 21]. The derived governing equations offer useful physical interpretations of the terms that arise in the projection process, and can be used to derive error bounds and consistency proofs for downstream reduced models.
A framework called quantum mechanical closure (QMCl) [55] has been recently developed which models the unresolved degrees of freedom as a quantum mechanical system. This approach can be viewed as a generalization of surrogate modeling based on stochastic models. Building on earlier work on quantum mechanical data assimilation [56, 57], QMCl encodes statistical information about the unresolved degrees of freedom in a quantum density operator [58, 59, 60] that evolves under the transfer operator of the dynamical system [61]. The fluxes of the unresolved variables are represented by quantum observables, while the generation of the required flux terms in the course of timestepping is carried out using tools from the mathematical theory of quantum measurement [58, 59, 60]. The resulting closure scheme is expressed in terms of linear operators acting on Hilbert spaces. This enables automatic positivity preservation (important for modeling sign definite physical quantities such as density or temperature) as well as data driven approximation using kernel integral operator methods [62, 63, 64] with rigorous consistency guarantees in the limit of infinite training data. QMCl has thus far been applied to the closure of prototype chaotic ordinary differential equations such as the Lorenz 63 and multiscale Lorenz 96 models, where it has shown promising results for the reconstruction of time correlation functions and long term statistics of observables.
1.1 Main results
In this work we extend the QMCl framework to the case of spatiotemporal dynamics governed by partial differential equations. In the extended framework, information about the unresolved degrees of freedom is encoded in fields of quantum density operators which are embedded within the spatial discretization cells of the coarse model. This produces a framework that can naturally take advantage of the spatial structure exhibited by partial differential equations, such as dynamical symmetries. More specifically, we employ elements of the vector valued spectral analysis (VSA) framework introduced in [65] to develop a data driven scheme that can effectively and efficiently model the closure of spatiotemporal dynamics in the presence of symmetries through the use of invariant basis functions. The closure framework can be used to derive both deterministic and stochastic surrogate models; in this work we focus on the deterministic setting.
The main features of the present work can be summarized as follows.
  1. 1.
    Positivity preservation. As with the original QMCl work, our framework uses the noncommutative setting of operator algebras to encode the original classical dynamics, leading to a closure scheme that is positivity preserving. This means that positive flux terms in the original dynamics are represented by positive operators in our closure model, producing positive surrogate flux terms. Generating surrogate fluxes that preserve the sign of the original fluxes is a nontrivial task and has important consequences for the closure of dynamics with sign definite quantities.
  2. 2.
    Symmetry factorization. Taking advantage of the symmetry factorization properties of the VSA framework, our dynamical closure scheme rests on a representation of the dynamics that is invariant under the actions of the spatial symmetry group of the resolved dynamics. This enables the formation of a compressed representation of the original spatiotemporal dynamics, leading to a data driven realization of our closure scheme that can be applied to complex dynamics with tractable computational cost.
  3. 3.
    Computational efficiency. In addition to the factorization of spatial symmetries, we implement the data driven formulation of our scheme in a way that lowers the cost of its most computationally expensive operations. As with the original QMCl and VSA frameworks, our data driven scheme requires solving an eigenvalue problem for a kernel matrix. The dimension of this matrix can be very large, making the eigenvalue problem one of the most expensive tasks in our scheme. To address that we compute a low rank approximation of the original kernel matrix and use it to solve the eigenvalue problem with reduced cost [66, 67].
  4. 4.
    Application. We apply our scheme to a closure problem for the shallow water equations on a one dimensional spatial domain. The shallow water equations are a hyperbolic system of partial differential equations with important applications in fields such as water waves modeling and ocean dynamics [68, 69]. The results demonstrate that our scheme can effectively capture and reproduce the main features of the dynamics, including for out of sample initial conditions.
1.2 Paper layout
In Section 2 we introduce the closure problem under study. This is followed by a brief overview of the original QMCl framework in Section 3. In Section 4 we extend this framework to the closure of spatiotemporal dynamics. We present the main features of the extended scheme, the elements it borrows from the VSA framework, as well as its data driven realization. In Section 5 we apply the extended scheme to a closure problem for the shallow water equations, present our numerical results and discuss some aspects of the closure scheme and future work directions. This is followed by a brief conclusion in Section 6. Finally, A provides additional details on the numerical implementation of the data driven formulation of our closure scheme.
2 Problem statement
@@
We consider a dynamical system $\Phi^{t}\colon X\rightarrow X$, $t\geq 0$ on a generally infinite dimensional state space $X$ possessing an invariant probability measure $\mu$. The flow map $\Phi^{t}$ represents the spatiotemporal dynamics governed by a partial differential equation on a bounded spatial domain $S\subset\mathbb{R}^{c}$ for some $c\in\mathbb{N}$. As such, elements of state space $X$ are functions on the domain $S$; more specifically, we consider $X\subset L^{2}(S,\nu;\mathbb{R}^{d})$ with $d\in\mathbb{N}$ and $\nu$ denoting the Lebesgue measure on $S$.
We decompose the state space as $X=\mathcal{X}\times\mathcal{Y}$ with $\mathcal{X}$ denoting the state space of the resolved degrees of freedom of the dynamics and $\mathcal{Y}$ the state space of the unresolved degrees. We assume that the resolved state space $\mathcal{X}\subset X$ is finite dimensional, a property that arises naturally in many applications. For instance many dissipative partial differential equations give rise to long time dynamics that is confined to a finite dimensional subset of their state space, such as an inertial manifold [70, 71]. In other cases discretization of the differential operator modeling the considered dynamics also yields a finite dimensional state space.
We define the Hilbert space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ with $H_{S}=L^{2}(S,\nu_{M};\mathbb{C})$ where $\nu_{M}$ denotes a restriction of the Lebesgue measure $\nu$ that will be constructed in Section 4.1. Every observable $\vec{f}\in H$ represents a spatiotemporal pattern of the dynamics, for which every state $x\in X$ yields a function on the spatial domain $\vec{f}(x)\in H_{S}$.
We introduce $d$ real observables $z_{j}\in H$, $j\in\{0,\ldots,d-1\}$ which will allow us to include in the resolved dynamics terms that depend on the unresolved variables. We refer to the observables $z_{j}$ as the flux functions and to their generated terms $z_{j}(x)\in H_{S}$ for any $x\in X$ as their fluxes. In what follows we use $z=\{z_{j}\}_{j=0}^{d-1}$ to denote the collection of flux functions. Based on our decomposition of state space $X$ we define the projected flow maps for the resolved dynamics $\Phi_{\mathcal{X}}^{t}\colon X\to\mathcal{X}$, $\Phi_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\Phi^{t}$ and unresolved dynamics $\Phi_{\mathcal{Y}}^{t}\colon X\to\mathcal{Y}$, $\Phi_{\mathcal{Y}}^{t}=\operatorname{proj}_{\mathcal{Y}}\circ\Phi^{t}$ where $\operatorname{proj}_{A}$ denotes the canonical projection onto a subspace $A\subset X$. With these definitions in place, the dynamics $\Phi^{t}$ can be expressed as
$\Phi^{t}(\hat{x},y)=\bigl{(}\Phi^{t}_{\mathcal{X}}(\hat{x},z(\hat{x},y)),\Phi_{\mathcal{Y}}^{t}(\hat{x},y)\bigr{)}$
with state vector $(\hat{x},y)\in\mathcal{X}\times\mathcal{Y}$.
Our goal in this work is to construct surrogate dynamics $\tilde{\Phi}^{t}\colon\tilde{X}\rightarrow\tilde{X}$ in a surrogate state space $\tilde{X}=\mathcal{X}\times\tilde{\mathcal{Y}}$ where $\tilde{\mathcal{Y}}$ denotes a surrogate unresolved state space. We introduce a surrogate version of the flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ and denote their collection by $\tilde{z}$. The surrogate dynamics $\tilde{\Phi}^{t}$ can now be expressed as
$\tilde{\Phi}^{t}(\hat{x},\tilde{y})=\bigl{(}\tilde{\Phi}_{\mathcal{X}}^{t}(\hat{x},\tilde{z}(\tilde{y})),\tilde{\Phi}_{\mathcal{Y}}^{t}(\hat{x},\tilde{y})\bigr{)}$ (1)
with state vector $(\hat{x},\tilde{y})\in\mathcal{X}\times\tilde{\mathcal{Y}}$, surrogate resolved dynamics $\tilde{\Phi}_{\mathcal{X}}^{t}\colon\tilde{X}\to\mathcal{X}$, $\tilde{\Phi}_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\tilde{\Phi}$ and surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{t}\colon\tilde{X}\to\tilde{\mathcal{Y}}$, $\tilde{\Phi}_{\mathcal{Y}}^{t}=\operatorname{proj}_{\tilde{\mathcal{Y}}}\circ\tilde{\Phi}$. To build the surrogate dynamics map $\tilde{\Phi}^{t}$ we need to define the surrogate space $\tilde{\mathcal{Y}}$, the surrogate flux functions $\tilde{z}_{j}$ and the surrogate dynamics $\tilde{\Phi}_{\mathcal{X}}^{t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{t}$.
@@
We are particularly interested in systems with dynamical symmetries. For this reason we assume that there exists a symmetry group $G$ acting on the spatial domain $S$ with a continuous nontrivial left action $\Gamma_{S,g}\colon S\rightarrow S$, $g\in G$ which preserves the null sets of measure $\nu$. The actions $\Gamma_{S,g}$ induce the actions on state space $\Gamma_{X,g}\colon X\rightarrow X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$, $g\in G$. The actions of group $G$ on $X$ represent dynamical symmetries of the dynamics $\Phi^{t}$; namely, we assume the following equivariance property
@@
for all $t\geq 0$ and $g\in G$. In such situations we seek to leverage the dynamical symmetries to reduce training data requirements and the overall dimension of our surrogate models.